OpenStudy (anonymous):

please help 5a = 5b -- -- x x-1

OpenStudy (princeharryyy):

wat does the ques mean ? wat do they want from the equations

OpenStudy (nurali):

(5a)/(x)= (5b)/(x-1) Multiplying by x(x-1) 5a(x-1)= 5bx Dividing by 5 a(x-1) = bx ax-a = bx ax-bx = a (keeping the x-terms on one side and the constant part on the other side) (changing sign while changing side) (a-b)x = a x = a/(a-b) (dividing by (a-b) ) Verification: (5a)/(x)= 5a/[a/(a-b)] = 5(a-b)(cancelling a) RHS = (5b)/(x-1) = 5b/[a/(a-b)-1] =5b/{[a-(a-b)]/(a-b)} =5b/{a-a+b/(a-b)} =5b/[b/(a-b)] =5(a-b) (cancelling b) =LHS

OpenStudy (anonymous):

 correct?

OpenStudy (nurali):

x = a/(a-b)

OpenStudy (anonymous):

ohhh ok gotcha!!